OpenStudy (anonymous):

Circle A has an equation of x2 + (y + 3)2 = 16. Where is the center of the circle located?

OpenStudy (anonymous):

what do you think is the center of this one?

OpenStudy (anonymous):

I would say, the X axis?

OpenStudy (anonymous):

(x )^2 + (y+3)^2 = 16 (x - h)^2 + (y+3)^2 = 16

OpenStudy (anonymous):

what is h in this case?

OpenStudy (anonymous):

x - ??? = x

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

x - 16 = x ??? that don't seem right does it? x - ??? = x for example, 7 - ??? = 7

OpenStudy (anonymous):

hmm.. im not sure.

OpenStudy (anonymous):

try this one: 10 - ? = 10 what do i subtract from 10 to get 10?

OpenStudy (anonymous):

still having a hard time?

OpenStudy (anonymous):

yea, sorry, I do not know why i don't know that :/

OpenStudy (anonymous):

it's ok... this is one part a lot of people mess up with too..

OpenStudy (anonymous):

say you have 20 dollars and you left it on the table... you think someone took away some money from the table and after you're count the amount of money on the table there is still 20 dollars. how much money was taken away?

OpenStudy (anonymous):

oh wow im so dumb(dim). haha. So it is 0?

OpenStudy (anonymous):

yes... and you're not dumb

OpenStudy (anonymous):

haha ty, i'm just joking. I was assessing the problem thinking it was something much more complicated when it was actually something simple as 0.

OpenStudy (anonymous):

yw... let's be clear though... what's the center of the circle?

OpenStudy (anonymous):

Would it be located in the 3rd quadrant? If not then i still don't get it :/

OpenStudy (anonymous):

y+3=0 -> y=-3 So the center is (0,-3) which is located right here. 

OpenStudy (anonymous):

yes ^^^, remember h=0 so your (h,k) is (0,-3)

OpenStudy (anonymous):

so what does y equal?